Let the enery stored "in the capacitor" be U. Show that U is given by the expression: U=½Q²/C


1
Expert's answer
2021-05-09T15:32:22-0400

A charged Capacitor is a store of electrical potential energy.

To find the energy stored in a capacitor, let us consider a capacitor of capacitance C, with a potential difference V between the plates.

There is a charge +q on one plate and –q on the other.

Suppose the capacitor is charged gradually. At any stage ,the charge on the capacitor is q.

Potential of capacitor "=\dfrac{q}{C}"


Small amount of work done in giving an additional charge "dq" to the capacitor is:


"dW=\dfrac{q}{C} .dq"


Integrating both sides,


"\int dW = \dfrac{1}{C}\int_{0}^{Q}qdq"


"W = \dfrac{Q^2}{2C}" "= U"